I am stuck on this problem.
(8/15 - 5/12) divide by 1/20.
I don't know what the first step is. Do i need to find the LCD, then what? I do know that I need to flip the 1/20.
Thank you for your help.
Laura

Answer by John10(297)

The first step is to simplify the numerator: 8/15 - 5/12
You do that by finding the LCD of 15 and 12 which is 60. Then you make each fraction has the same denominator of 60 and subtract the numerator.
8/15 - 5/12 = 32/60 - 25/60 = 7/60
Next you will divide 7/60 by 1/20 . You do that by multiplying 7/60 by 20
(7/60) * (20) = 7/3
Hope it helps you:)
John10